two points b and c lie on radial line of a rotating disk. the points are 2 inches apart. vb=700 ft/min and vc=880 ft/min. find the radius of rotation for each of these points. slader
1
Expert's answer
2020-09-29T14:10:11-0400

Let the radius be r

let us assume C is at the edge

so distance from center"= r ft"

B will then be = "(r - \frac{2}{12}) ft"

angular velocity will be same hence

"\frac{Vb}{Rb}= \frac{Vc}{Rc}"


"\frac{700}{(r - \frac{1}{6})} = \frac{880}{r}"


"\frac{r}{(r - \frac{1}{6})} = 1.257"


"r = 1.257r - 0.2095"


"0.257r = 0.2095"


"r = 0.815 ft"


Hence C is at 0.815 ft and B is at 0.648 ft